Find an equation of the parabola passing through the points (0,4), (1,2), and (-1,8)
Answer by lwsshak3(11628)

Find an equation of the parabola passing through the points (0,4), (1,2), and (-1,8)
**
Quadratic equation:
y=Ax^2+Bx+C
..
(0,4)
4=0+0+C
c=4
..
(1,2)
2=A+B+4
..
(-1,8)
8=A-B+4
..
2=A+B+4
8=A-B+4
add
10=2A+8
2A=2
A=1
B=2-A-4=-3
C=4
..
Equation of parabola:
y=x^2-3x+4